# Watch and Learn: Learning to Use Computers from Online Videos

### v1

@@ Abstract
Computer use agents (CUAs) need to plan task workflows grounded in diverse, ever-changing applications and environments, but learning is hindered by the scarcity of large-scale, high-quality training data in the target application. Existing datasets are domain-specific, static, and costly to annotate, while current synthetic data generation methods often yield simplistic or misaligned task demonstrations. To address these limitations, we introduce Watch & Learn (W&L), a framework that converts human demonstration videos readily available on the Internet into executable UI trajectories at scale. Instead of directly generating trajectories or relying on ad hoc reasoning heuristics, we cast the problem as an inverse dynamics objective: predicting the user’s action from consecutive screen states. This formulation reduces manual engineering, is easier to learn, and generalizes more robustly across applications. Concretely, we develop an inverse dynamics labeling pipeline with task-aware video retrieval, generate over 53k high-quality trajectories from raw web videos, and demonstrate that these trajectories improve CUAs both as in-context demonstrations and as supervised training data. On the challenging OSWorld benchmark, UI trajectories extracted with W&L consistently enhance both general-purpose and state-of-the-art frameworks in-context, and deliver stronger gains for open-source models under supervised training. These results highlight web-scale human demonstration videos as a practical and scalable foundation for advancing CUAs towards real-world deployment.
Figure 1: W&L converts web-scale human demonstration videos into executable UI trajectories, providing scalable supervision and in-context exemplars for computer use agents.
@@ 1 Introduction
Computer use agents (CUAs) (Zheng et al., 2024a; Kil et al., 2024; Qin et al., 2025; Gou et al., 2025; OpenAI, 2025b) hold the promise of transforming how humans interact with software and the web, from everyday productivity tasks to enterprise-scale automation. To be effective, CUAs must both plan multi-step task workflows that incorporate domain knowledge, and ground these plans into concrete UI actions within diverse and ever-changing applications. Progress toward these capabilities hinges on access to high-quality task demonstrations, yet collecting annotated trajectories at scale is prohibitively expensive.
Meanwhile, the web is rich in human demonstration videos (e.g., YouTube tutorials, screencasts, etc.), which naturally encode complex workflows across diverse applications. Unlocking this resource could provide CUAs with scalable supervision and rich priors for expert-level planning. However, existing synthetic data generation approaches have fallen short of realizing this vision.
@@
Prior efforts fall into three main categories: Offline synthesis attempts to recover trajectories from videos using pipelines that combine multimodal large language models (MLLMs) with UI element detectors and transition parsers. Despite substantial engineering, systems such as MONDAY (Jang et al., 2025b) and TongUI (Zhang et al., 2025) achieve only modest action labeling accuracies ($\sim$70% for MONDAY), reflecting the limitations of multi-stage heuristics. Online synthesis generates trajectories through random exploration in real-world environments and later retrofits them with pertinent task instructions (Murty et al., 2024; Sun et al., 2025). While scalable in principle, this approach produces low-complexity demonstrations that are less aligned with human goals and can be costly as they require online exploration. Hybrid approaches, such as Explorer (Pahuja et al., 2025), generate task proposals and then execute and refine them online, but still rely on MLLMs for action grounding—thereby sharing similar limitations to offline synthesis methods.
Overall, these approaches either rely on brittle heuristics, are costly as they rely on explorations in real environments, or generate low-complexity demonstrations misaligned with human intent. To address these limitations, this work introduces Watch & Learn (W&L), a framework that converts human demonstration videos readily available online into executable UI trajectories at scale (Figure 1). Instead of directly generating trajectories or depending on complex multi-stage pipelines, we frame the problem as an inverse dynamics objective: given two consecutive observations $(O_{t},O_{t+1})$, predict the intermediate action $a_{t}$ that produced the transition. This formulation is easier to learn, avoids hand-crafted heuristics, and generalizes robustly across applications. In robotics, inverse dynamics modeling is a well-established method for recovering actions from state transitions (e.g., VPT (Baker et al., 2022), DreamGen (Jang et al., 2025a)); here, we demonstrate that the same principle can be adapted effectively for CUAs. From our experiments, this simple formulation yields a highly accurate model of user behavior, sidestepping the complexity of conventional pipelines.
To scale this approach to the web, we construct a large state-transition corpus of 500k state transition data from real-world web interactions. Each example consists of an observation at time $t$, an action, and the resulting observation at $t+1$. Training an inverse dynamics model (IDM) on this corpus allows us to directly map visual transitions into structured actions. We further design a retrieval framework that retrieves YouTube videos relevant to target tasks (for in-context learning) or general video tutorials (for supervised fine-tuning). Applying the IDM to these videos transforms raw demonstrations into high-quality trajectories, covering a broad spectrum of real-world workflows.
Beyond data collection, W&L uncovers a different role for CUAs. In addition to effectively using UI trajectories in training, we demonstrate that the extracted trajectories can also serve as in-context exemplars during inference, enabling CUAs to leverage planning and grounding priors enriched with domain knowledge on the fly. This dual role (training and in-context guidance) enables flexible integration with both open-source models and general-purpose agents. To illustrate the effectiveness of this approach, we evaluate W&L on OSWorld (Xie et al., 2024), a challenging benchmark requiring both domain familiarity and strong planning and grounding capabilities. On OSWorld, trajectories extracted from web-scale videos deliver consistent gains: in-context use improves general-purpose models and state-of-the-art agentic frameworks by up to 3 percentage points, while training with them yields even larger improvements for open-weight models (up to 11 percentage points). Importantly, these benefits are achieved without any manual annotation, demonstrating that web-scale human workflows can serve as a practical and scalable foundation for advancing CUAs towards real-world deployment.
In summary, our contributions are three-fold: (i) We develop a scalable inverse dynamics labeling pipeline, coupled with a task-aware video retrieval framework, that transforms raw web videos into high-quality trajectories. Overall, without any manual effort, we generate 53,125 trajectories with high-accuracy action labels. (ii) We show that these video-derived trajectories can serve as in-context demonstrations at inference time, improving general-purpose CUAs without retraining. (iii) We also demonstrate that these trajectories provide effective training data, offering a scalable supervision signal that substantially improves open-source CUAs.
@@ 2 Related Work
2.1 Data Synthesis for Computer Use Agents
While human-curated UI control datasets have been collected (Deng et al., 2023; Lù et al., 2024; Rawles et al., 2023; Li et al., 2024), their limited size and diversity remains a key bottleneck for CUAs. Recent work has focused on synthesizing data from exploration, tutorials, or self-play.
Exploration-based approaches such as BAGEL (Murty et al., 2024), NNetNav (Murty et al., 2025), Explorer (Pahuja et al., 2025), and OS-Genesis (Sun et al., 2025) generate training data by letting agents explore websites and retroactively labeling their interactions with task instructions. This paradigm yields scalable but often noisy data, with alignment and accuracy depending heavily on heuristics or MLLM labeling. Other methods leverage online resources: Synatra (Ou et al., 2024) and AgentTrek (Xu et al., 2025) transform textual tutorials into executable trajectories, while TongUI (Zhang et al., 2025) aggregates a massive corpus of multimodal tutorials (text and screencast videos) into GUI interaction data. These approaches demonstrate that web-scale instructional content can provide diverse coverage across applications, but they rely primarily on off-the-shelf MLLMs to label trajectories, which often introduces brittleness or misalignment.
Another line of work integrates synthesis into the training loop itself. OpenWebVoyager (He et al., 2025) improves through online exploration and feedback; WebRL (Qi et al., 2025) generates new instructions from failed tasks to form a self-evolving curriculum; SCA (Qi et al., 2025) has agents self-generate and verify new tasks in a code-as-task format; and ZeroGUI (Yang et al., 2025) proposes a fully automated online learning framework for GUI agents, where VLMs generate tasks and rewards that drive reinforcement learning without manual annotations. These strategies enable continual improvement without additional human data, but often produce simplistic or narrow task distributions. Moreover, the process can be expensive as it involves multiple iterations of data generation and training.
Our framework, Watch & Learn, also leverages web videos like TongUI (Zhang et al., 2025), but differs in its technical strategy. Instead of relying on MLLMs to label tutorial steps, we train an inverse dynamics model (IDM) that can accurately infer user actions from consecutive screen states. This produces highly reliable UI trajectories that not only provide stronger supervised training signals but also serve as more effective in-context exemplars at inference time. By combining web-scale video mining with accurate action labeling, our approach complements prior work and highlights the value of extracting accurate cues from video-based supervision for CUAs.
@@ 2.2 In-context Learning for Agents
Despite being generally helpful, the effectiveness of ICL is heavily influenced by the scale of the LLMs and the size of their context window, particularly for long-horizon, multi-step tasks. While including more ICL examples usually brings performance gains (Agarwal et al., 2024), this method incurs significant computational overhead and latency with long demonstration trajectories. Therefore, efficiently selecting demonstration sequences (Gupta et al., 2025) or abstracting them in high-level workflows (Wang et al., 2024; Zheng et al., 2024b) has become a promising research direction. For computer-use agents, where tasks are often long and complex, one major challenge is the model’s inability to plan effectively. Several pieces of work have leveraged ICL to address this specific problem (Holt et al., 2025; Zhao et al., 2025).
@@ 3.1 Inverse Dynamics Model
A key component of our framework is an IDM that predicts the user action given two consecutive screen observations. Training such a model requires large-scale state-transition data, which is scarce in existing datasets. To address this gap, we construct our own corpus of transitions by synthesizing interactions at scale, complemented by existing human-collected datasets.
State-transition data collection. To obtain large-scale supervision, we built an automated data generation pipeline that interacts with live web pages and records state transitions. Inspired by WebDreamer (Gu et al., 2025), we randomly select entry points from the March 2025 Common Crawl index and launch browsing sessions that perform sequences of actions such as clicking, typing text, scrolling, and moving the cursor. The action policy is not uniform: we weight the sampling toward common interactions (e.g., clicks) while still ensuring that less frequent actions are covered. Through this procedure, we collected around 500k synthetic transitions. To complement these, we also incorporate 132k human-annotated transitions from the Mind2Web dataset (Deng et al., 2023), yielding a training corpus of more than 630k $(O_{t},a_{t},O_{t+1})$ triples.
@@
Model architecture. The IDM takes as input two consecutive observations $(O_{t},O_{t+1})$ and outputs the action $a_{t}$ that caused the transition. We adopt a vision-only architecture consisting of a SigLIP-2 vision encoder followed by four Transformer (Vaswani et al., 2017) layers. On top of this backbone, we attach three specialized prediction heads:
@@
  • Action classification head: a categorical predictor over five supported primitives: click, scroll, type, wait, and move.
@@
  • Coordinate head: for location-based actions (click, move, type), the model predicts normalized $(x,y)$ coordinates discretized into integers from 0 to 1000. This converts coordinate regression into a classification problem, which proved to be more stable in training.
  • Language head: for text entry actions, the model generates the string input using a GPT-2 small decoder (Radford et al., 2019) attached to the Transformer backbone.
@@
Scroll and wait actions require no additional arguments; the model simply predicts their occurrence.
Training and evaluation. The IDM is trained with a multi-task objective: cross-entropy for action class prediction, cross-entropy for discretized coordinates, and language modeling loss for text generation. Training is performed end-to-end over the 630k transition corpus. We evaluate the IDM on the held-out test split of Mind2Web (Deng et al., 2023), which provides human-annotated trajectories across diverse websites. This benchmark allows us to measure both action classification accuracy and argument prediction quality in a realistic setting. As reported in Section 4.2.2, our IDM trained on state transition data achieves stronger action accuracy than off-the-shelf foundation models, validating its effectiveness as the core labeling module in our framework.
3.2 Data Generation from Videos
@@
Once the IDM is trained, we retrieve suitable tutorial videos and apply the IDM.
@@
Video retrieval. We build a retrieval framework that searches and downloads tutorial videos from large video platforms such as YouTube. The retrieval procedure differs depending on whether the goal is inference-time support or large-scale training data collection. Inference-time retrieval. Given a task description and the target application, we form a natural language search query. To refine the query, we prompt Gemini 2.5 Flash [^1] https://generativelanguage.googleapis.com/v1beta/models/gemini-2.5-flash:generateContent (Gemini Team, 2025) with both the task instruction and the initial screen, asking it to generate a more specific query. We then use the YouTube Search API to retrieve the top 15 videos. For example, a task instruction "Can you increase the max volume of the video to the 200% of the original volume in VLC?" becomes the search query "vlc increase max volume". Each retrieved video is paired with its title, which we treat as the candidate task description. Training-time retrieval. To construct a broad training dataset, we curate a list of 69 applications spanning productivity, programming, design, screen editing, audio production, system utilities, and science/data domains. For each one, we prompt Gemini 2.5 Flash to generate plausible task queries and use them to search on video platforms, downloading the corresponding tutorial videos.
@@
Filtering. Not all retrieved videos are usable. We sample frames at 1 frame per second and automatically filter out segments that are not screencasts (e.g., talking-head segments), are zoomed in/out, or are blurred due to transitions. Gemini 2.5 Flash is used as a classifier to perform this filtering. For inference-time retrieval, we retain only the top 3 videos that pass filtering to minimize noise. For training data collection, we keep all videos that satisfy the filter.
@@
3.3 Applications of Trajectories
The trajectories extracted from videos can be used in two complementary ways: as in-context exemplars that guide models at inference time, and as supervised data that improve models via fine-tuning.
3.3.1 In-Context Learning
For in-context learning (ICL), we transform each trajectory into a demonstration that can be inserted directly into a model’s context window. Each trajectory consists of (observation, action) pairs, but simply showing raw frames and actions may not provide sufficient signal. To improve performance, we prompt Gemini 2.5 Flash to generate natural language rationales for each action in the trajectory, yielding demonstrations of the form $(\text{observation},\text{action},\text{reasoning})$. We format a small set of such demonstrations (typically 3–5) into the input prompt of a general-purpose agent model. At inference time, the agent is conditioned on these exemplars when predicting the next action for a new task, allowing it to draw on planning and grounding priors as well as application-specific knowledge distilled from real demonstrations, without additional training.
3.3.2 Supervised Fine-Tuning
For supervised fine-tuning (SFT), we aggregate the automatically labeled trajectories into a large-scale training corpus. Each trajectory is represented as a sequence of (state,action) pairs and used to optimize a multimodal large language model with a standard sequence modeling objective. We train two distinct model families. First, we fine-tune UI-TARS-1.5 (Qin et al., 2025), a strong, open source vision-language-action model designed specifically for computer use. This setting tests whether our trajectories can improve a model that already incorporates domain-specific priors. Second, we fine-tune Qwen 2.5-VL (Bai et al., 2025), a state-of-the-art open-weight multimodal LLM. This setting evaluates whether our data can also benefit general-purpose multimodal models that are not tailored to computer use. Overall, these experiments demonstrate our data’s value as a versatile supervision signal, capable of enhancing both specialized CUAs and large, open-source MLLMs.
4 Experiments
@@
4.1 Setup
4.1.1 Models
We evaluate three classes of models.
General-purpose multimodal models. Gemini 2.5 Flash (Gemini Team, 2025), OpenAI o3 (OpenAI, 2025a), and Claude 4 Sonnet (Anthropic, 2025) are tested in the in-context learning setting.
Agentic framework. We use Jedi (Xie et al., 2025), a state-of-the-art vision-only agentic framework for OSWorld. Jedi couples an MLLM planner (OpenAI o3), which outputs natural-language action steps, with the Jedi-7B grounding model, which maps those steps to executable UI actions. We report results both with and without our trajectories provided as in-context exemplars to the agent.
Open-source models. We train UI-TARS-1.5-7B (Qin et al., 2025) and Qwen 2.5-VL 7B (Bai et al., 2025) with supervised fine-tuning on our 53,125 video-derived trajectories. This dual evaluation highlights that our data improve both specialized CUAs and general-purpose multimodal models.
4.1.2 Datasets
Our experiments involve three categories of data.
State-transition corpus. To train the IDM, we collect approximately 500k transitions from autonomous web interactions and add 132k human-annotated transitions from Mind2Web (Deng et al., 2023), resulting in over 630k $(O_{t},a_{t},O_{t+1})$ triples.
Video-derived trajectories. Once trained, the IDM is applied to retrieved and filtered YouTube tutorials, producing 53,125 high-quality trajectories across 69 applications spanning productivity, programming, design, screen editing, audio production, system utilities, and scientific/data domains. The category distribution of these trajectories is summarized in Table 1.
As a data labeling baseline, we use TongUI (Zhang et al., 2025), which generates action annotations by prompting the UI-TARS-7B agent. Unlike our video-derived trajectories, these labels are often noisy and inaccurate due to reliance on an imperfect web agent, but they serve as a useful point of comparison for evaluating label quality.
Evaluation benchmark. We use OSWorld-Verified (Xie et al., 2024), the most up-to-date version of OSWorld, as our primary benchmark. It evaluates agents in real desktop and operating system environments across productivity, programming, design, and system utilities. Tasks must be solved under interactive execution with a 50-step limit, stressing agents’ ability to plan, ground instructions in dynamic states, and apply domain knowledge across diverse applications. This makes OSWorld-Verified a comprehensive testbed for both in-context learning and supervised fine-tuning.
@@ 4.1.2 Datasets
w/ video; Labeling: TongUI 21.1 (-0.7)
w/ video; IDM: W&L 24.3 (+2.5)

### v2

@@ Abstract
Computer-using agents (CUAs) must plan task workflows across diverse and evolving applications, yet progress is limited by the lack of large-scale, high-quality training data. Existing datasets are narrow, static, and costly to annotate, while synthetic data often yields oversimplified or misaligned behaviors. We present Watch & Learn (W&L), a framework that converts readily available Internet videos of human computer use into executable UI trajectories at scale. Instead of directly generating actions or relying on handcrafted heuristics, we cast trajectory annotation as an inverse dynamics problem that predicts user actions from consecutive screen states, which simplifies learning and generalizes across domains. Through a task-aware retrieval and labeling pipeline, W&L yields over 53K high-quality trajectories that enhance CUAs both as in-context exemplars and as supervised training data. On OSWorld, it consistently improves general-purpose and specialized CUAs, while on WindowsAgentArena it achieves state-of-the-art performance among 7B-scale models under the 15-step limit. These results show that web-scale human demonstration videos can serve as a practical and scalable foundation for advancing real-world CUAs.
Figure 1: W&L converts web-scale human demonstration videos into executable UI trajectories, providing scalable supervision and in-context exemplars for computer-using agents.
@@ 1 Introduction
Computer-using agents (CUAs) (Zheng et al., 2024a; Kil et al., 2024; Qin et al., 2025; Gou et al., 2025; OpenAI, 2025b; Yang et al., 2025b; Wang et al., 2025; Xie et al., 2025; Agashe et al., 2025a, b; Gonzalez-Pumariega et al., 2025) are emerging as a powerful paradigm for automating software and web interactions. From everyday productivity tools to enterprise workflows, CUAs must both plan multi-step task sequences that incorporate domain knowledge and ground these plans into concrete UI actions across diverse, constantly changing interfaces. Despite rapid progress, their scalability remains constrained by the difficulty of obtaining large-scale, high-quality demonstrations of human-computer interaction (Wang et al., 2025).
Collecting such trajectories remains extremely costly and time-consuming. For instance, OpenCUA’s AgentNet dataset (Wang et al., 2025), which spans three operating systems and more than 200 applications and websites, required six months of dedicated annotation effort to label 22K tasks, with a total cost exceeding \$32,000 (about \$1.45 per task). Despite its impressive coverage, AgentNet underscores the scalability bottleneck of manual annotation: extending to the millions of trajectories needed for robust generalization would require substantial additional time and cost well over half a million dollars, even for well-funded labs. Such high costs and long turnaround times make it difficult to build datasets that capture the full diversity and complexity of real-world computer use, limiting progress in developing scalable computer-using agents.
Meanwhile, the web already contains an enormous volume of human demonstration videos, including tutorials, screencasts, and walkthroughs, which naturally encode rich task workflows across applications. If we could reliably convert these videos into structured action trajectories, CUAs could learn directly from human expertise at unprecedented scale. However, existing trajectory generation methods have struggled to achieve this vision.
@@
Prior work falls into three broad categories. Offline synthesis reconstructs trajectories from recorded videos using multi-stage pipelines that couple multimodal large language models (MLLMs) with UI element detectors and transition parsers. Despite substantial engineering, systems such as MONDAY (Jang et al., 2025b) and TongUI (Zhang et al., 2025) report modest accuracies (around 70% for MONDAY) due to cascading errors in heuristic pipelines. Online synthesis collects trajectories through random exploration and retrofits them with task instructions (Murty et al., 2024; Sun et al., 2025), but these data are low in task complexity and expensive to gather. Hybrid approaches, such as Explorer (Pahuja et al., 2025), mix proposal generation with online refinement, yet still depend on MLLMs for action grounding, retaining many of the same limitations.
To address these limitations, we introduce Watch & Learn (W&L), a framework that formulates trajectory recovery as an inverse dynamics prediction problem. Given two consecutive screen observations $(O_{t},O_{t+1})$, the model predicts the intermediate action $a_{t}$ that produced the transition. This one-step formulation is substantially easier to learn than generating entire trajectories, avoids hand-crafted parsing heuristics, and generalizes across applications. In robotics, inverse dynamics modeling has long been used to infer actions from state transitions (e.g., VPT (Baker et al., 2022), DreamGen (Jang et al., 2025a)); here, we adapt this principle to computer use, showing it can recover accurate, executable actions from raw visual transitions.
We train an inverse dynamics model (IDM) on a large state-transition corpus containing more than 600K automatically labeled transitions from real-world web interactions. Each sample records $(O_{t},a_{t},O_{t+1})$, enabling direct supervision from interaction logs without manual annotation. We then pair this model with a task-aware retrieval module that locates relevant YouTube videos, either those specific to target benchmarks or general instructional content, and automatically converts them into structured UI trajectories. This pipeline yields 53,125 high-quality trajectories spanning a broad range of applications and task categories across three operating systems (Ubuntu, macOS, and Windows).
Beyond scalable data collection, W&L reveals a dual role for such trajectories. They can be used not only for supervised fine-tuning, but also as in-context exemplars at inference time, allowing CUAs to dynamically retrieve demonstrations as planning and grounding priors. This enables flexible integration with both open-source CUAs and general-purpose multimodal agents.
We evaluate W&L on two challenging benchmarks, OSWorld (Xie et al., 2024) and WindowsAgentArena (Bonatti et al., 2024), which require complex reasoning and precise UI grounding. On OSWorld, trajectories extracted from web videos yield consistent gains: in-context use improves general-purpose models and agentic frameworks by up to 3 percentage points, while supervised fine-tuning provides larger boosts for open-weight models (up to 11 points). On WindowsAgentArena, fine-tuning with our trajectories achieves leading performance among 7B-scale models under the 15-step evaluation limit, demonstrating that large-scale web demonstrations can translate effectively into high-quality training data. These benefits are obtained without any manual annotation, highlighting the potential of web-scale human workflows as a practical and scalable foundation for advancing CUAs toward real-world deployment. In summary, this work makes the following contributions:
  • We introduce a inverse dynamics model that recovers UI actions directly from visual transitions, removing the need for heuristic parsing or manual labeling.
  • We develop a task-aware retrieval framework that converts web-scale human demonstration videos into 53K executable UI trajectories.
  • We show that these trajectories enhance CUAs both as in-context exemplars and as supervised training data, achieving strong gains on OSWorld and WindowsAgentArena.
@@ 2 Related Work
2.1 Data Synthesis for Computer-Using Agents
Human-curated UI datasets (Deng et al., 2023; Lù et al., 2024; Rawles et al., 2023; Li et al., 2024) remain limited in scale and diversity, constraining progress in CUAs. Recent research therefore explores synthesizing data through exploration, tutorials, or self-play.
Exploration-based synthesis methods such as BAGEL (Murty et al., 2024), NNetNav (Murty et al., 2025), Explorer (Pahuja et al., 2025), and OS-Genesis (Sun et al., 2025) let agents explore websites and retroactively label interactions with task instructions. This yields large-scale data but often introduces noise, since quality depends heavily on heuristics or multimodal LLM labeling.
Tutorial-driven synthesis leverages online instructional resources. Synatra (Ou et al., 2024) and AgentTrek (Xu et al., 2025) convert textual tutorials into executable trajectories, while TongUI (Zhang et al., 2025) aggregates multimodal tutorials (text and screencast videos) into GUI data. These works show that web-scale content can broaden coverage across applications, though reliance on off-the-shelf LLMs often leads to brittleness and misalignment.
Self-improving agents integrate data synthesis directly into training. OpenWebVoyager (He et al., 2025) improves via online exploration and feedback, WebRL (Qi et al., 2025) creates new instructions from failures, SCA (Qi et al., 2025) generates and verifies code-based tasks, and ZeroGUI (Yang et al., 2025a) trains GUI agents entirely through automated task and reward generation. These systems enable continual learning without human data but often yield narrow or repetitive task distributions and require expensive iterative retraining.
Our approach, Watch & Learn, also mines web videos like TongUI but differs fundamentally in supervision. Rather than labeling videos with MLLMs, we train an inverse dynamics model that infers user actions directly from screen transitions. This produces accurate UI trajectories that strengthen both supervised training and in-context reasoning. By combining large-scale video mining with precise action labeling, our framework complements prior synthesis efforts and emphasizes the value of reliable video-based supervision for CUAs.
@@ 2.2 In-context Learning for Agents
Despite being generally helpful, the effectiveness of ICL is heavily influenced by the scale of the LLMs and the size of their context window, particularly for long-horizon, multi-step tasks. While including more ICL examples usually brings performance gains (Agarwal et al., 2024), this method incurs significant computational overhead and latency with long demonstration trajectories. Therefore, efficiently selecting demonstration sequences (Gupta et al., 2025) or abstracting them in high-level workflows (Wang et al., 2024; Zheng et al., 2024b) has become a promising research direction. For computer-using agents, where tasks are often long and complex, one major challenge is the model’s inability to plan effectively. Several pieces of work have leveraged ICL to address this specific problem (Holt et al., 2025; Zhao et al., 2025; Agashe et al., 2025b, a; Gonzalez-Pumariega et al., 2025).
@@ 3.1 Inverse Dynamics Model
A key component of our framework is an inverse dynamics model (IDM) that predicts the user action given two consecutive screen observations. Since large-scale state-transition data are scarce, we construct a new corpus by synthesizing interactions at scale.
State-transition data. To obtain large-scale supervision, we built an automated pipeline that interacts with live web pages and records transitions. Inspired by WebDreamer (Gu et al., 2025), we sample entry points from the March 2025 Common Crawl index and launch browsing sessions that execute actions drawn from the above schema. The sampling policy favors frequent actions (e.g., click) while ensuring sufficient diversity. This process yields roughly 600K $(O_{t},a_{t},O_{t+1})$ transitions. Further implementation details are provided in the Appendix.
Action space. The IDM models six atomic primitives that together capture the majority of real-world computer interactions: Click, Release, Scroll, Type, Wait, and Move. Each action is parameterized by coordinates, text, or scalar values as appropriate. The Click primitive corresponds to a press event and includes normalized screen coordinates $(x,y)$, whereas Release denotes the corresponding button-up event and requires no positional arguments. This separation simplifies supervision and aligns with low-level GUI event semantics, where drag-and-drop gestures can be represented compositionally as a sequence of Click $\rightarrow$ Move $\rightarrow$ Release. Modeling these events as distinct atomic classes allows the IDM to predict a single action per transition while maintaining the flexibility to represent both short and prolonged clicks. Overall, this action space balances coverage and learnability, providing sufficient expressiveness for all interaction types required by OS-level benchmarks such as OSWorld and WindowsAgentArena.
@@
Model architecture. The IDM takes as input two consecutive observations $(O_{t},O_{t+1})$ and outputs the action $a_{t}$ that caused the transition. Inspired by IDM architectures for robotics (Baker et al., 2022; Jang et al., 2025a), we adopt a vision-only architecture consisting of a SigLIP-2 vision encoder (Tschannen et al., 2025) followed by four Transformer (Vaswani et al., 2017) layers. On top of this backbone, we attach three specialized prediction heads:
@@
  • Action classification head: a categorical predictor over six supported primitives: click, release, scroll, type, wait, and move.
@@
  • Coordinate head: for location-based actions (click, move, and type), the model predicts normalized $(x,y)$ coordinates discretized into integers from 0 to $999$. This formulation converts coordinate regression into a classification problem, which proved more stable during training.
  • Language head: for text-entry actions, the model predicts the typed string autoregressively using a lightweight GPT-2 decoder (Radford et al., 2019) conditioned on the visual features of $(O_{t},O_{t+1})$. This design enables variable-length text generation while maintaining grounding in the observed state transition.
@@
Scroll, wait, and release actions require no additional arguments; the model simply predicts their occurrence.
Training. The IDM predicts three types of outputs depending on the action: (1) an action primitive, (2) discretized pixel coordinates $(x,y)$ for location-based actions, and (3) token sequences for text-entry. Given a transition $(O_{t},O_{t+1})$, the model outputs a distribution over action primitives, optional coordinate distributions over $1000$ bins per axis (integers 0999999), and an autoregressive distribution over text tokens. In practice, we discretize coordinates instead of regressing in continuous space, which stabilizes training and yields more accurate click and typing locations.
We train the model with a multi-task objective that activates only the losses relevant to the ground-truth action. Concretely, we apply a classification loss for the primitive, a coordinate loss for location-based actions, and a language-modeling loss for text-entry. Each head uses a standard cross-entropy term, and all losses are weighted equally unless otherwise specified. Full loss definitions are in the Appendix.
Evaluation. We evaluate the IDM on a held-out subset of the state-transition corpus containing 100 transitions per action type. This evaluation set spans diverse real-world web interactions collected using the same automated pipeline described earlier. We assess performance along two dimensions: ActionType accuracy, which measures the correctness of the predicted action primitive, and Action accuracy, which additionally includes the precision of the predicted coordinates and generated text. As shown in Section 4.4.1, the IDM trained solely on state-transition supervision achieves strong action classification accuracy, outperforming off-the-shelf foundation models and confirming its effectiveness as the labeling backbone of our framework.
3.2 Trajectory Data Generation from Videos
@@
Once the IDM is trained, we retrieve suitable tutorial videos and use the IDM to convert them into UI trajectories.
@@
Video retrieval. We build a retrieval framework that searches and downloads tutorial videos from large video platforms such as YouTube. The retrieval procedure differs depending on whether the goal is inference-time support or large-scale training data collection. Inference-time retrieval. Given a task description and the target application, we form a natural language search query. To refine the query, we prompt Gemini 2.5 Flash [^1] https://generativelanguage.googleapis.com/v1beta/models/gemini-2.5-flash:generateContent (Gemini Team, 2025) with both the task instruction and the initial screen, asking it to generate a more specific query. We then use the YouTube Search API to retrieve the top 15 videos. For example, a task instruction "Can you increase the max volume of the video to the 200% of the original volume in VLC?" becomes the search query "vlc increase max volume". Each retrieved video is paired with its title, which we treat as the candidate task description. Training-time retrieval. To construct a broad training dataset, we curate a list of 69 applications spanning productivity, programming, design, screen editing, audio production, system utilities, and science/data domains (Table 1). For each application, we prompt Gemini 2.5 Flash to generate plausible task queries and use them to search on video platforms, downloading the corresponding tutorial videos.
@@
Filtering. Retrieved videos vary widely in quality and content, so we first sample frames at 1 frame per second and discard unusable segments. Segments that are not screencasts (e.g., talking-head clips), are zoomed or cropped, or are blurred due to scene transitions are automatically removed by prompting Gemini 2.5 Flash. At inference time, we retain only the top 3 videos that pass this filter to reduce noise, whereas for training data collection we keep all qualified videos.
@@
Table 2: Main results on OSWorld-Verified (Xie et al., 2024) under the 50-step evaluation limit. IDM-labeled video exemplars from W&L improve both general multimodal models (Gemini, Claude, o3, Jedi) via ICL and open-source CUAs via SFT, achieving up to +11.1+11.1 absolute gains over TongUI-labeled and baseline models.
@@
w/ TongUI labeled videos 21.1 (-0.7)
w/ W&L labeled videos 24.3 (+2.5)